a star can become a black hole if all of its mass is within its schwarzchil radius . the sun which has a mass of 1.99*10^30kg , will eventually burn out and collapse . what would be its schawarzchilds radius , if it was possible for it to be a black hole ?
Answer #1

we know, escape speed, ve = sqrt(2*G*M/R)

when, R = Rs( schwarzchil radius) ==> Ve = c

c = sqrt(2*G*M/Rs)

c^2 = 2*G*M/Rs

==> Rs = 2*G*M/c^2

= 2*6.67*10^-11*1.99*10^30/(3*10^8)^2

= 2950 m or 2.95 km